Hi, can anyone help me figure out this co-ordinate geometry problem? I am having a tough time with it...
Find the equation of the line passing through the mid-point of (2,4) and (-6,-2) and also passing through (8,7). Leave answer in general form.
Thank you! I have a test on these tomorrow.

Answer by Gogonati(855)

The midpoint of the line segment is: (2-6)/2=-2 and (4-2)/2=1, (-2,1)
The slope of the line that pases at points (-2,1) and (8,7) is m=(1-7)/(-2-8)=3/5
The equation of the line with slope 3/5 and point (8,7) is: y-7=3/5(x-8) or
5(y-7)=3(x-8), 3x-5y=-11.